OpenStudy (anonymous):

Find a parabola, passing through the origin, whose tangent line at (2,4) passes through the point (0,8).

OpenStudy (anonymous):

ok if it passes through the origin you know that \[f(x)=ax^2+bx=c\] and \[f(0)=c=0\] so right away you know the constant is 0

OpenStudy (anonymous):

otherwise it cannot pass through the origin. constant is 0. now you know that the tangent line at the point (2,4) passes through (0,8)

OpenStudy (anonymous):

the slope of that line is \[\frac{8-4}{0-2}=\frac{4}{-2}=-2\]

OpenStudy (anonymous):

this means that the derivative at the point (2,4) is -2

OpenStudy (anonymous):

\[f(x)=ax^2+bx\] \[f'(x)=2ax+b\] \[f'(2)=4a+b=-2\] so we know \[4a+b=-2\]

OpenStudy (anonymous):

and finally we know that (2,4) is on the graph of f, which means \[f(2)=a\times 2^2+b\times 2 = 4\] \[4a+2b=4\]

OpenStudy (anonymous):

so we have two equations and two unknowns, namely \[4a+b=-2\] and \[4a+2b=4\] subtract first from second to get b=6 etc

OpenStudy (anonymous):

\[4a+6=-2\] \[4a=-8\] \[a=-2\] if my arithmetic is write and your function is \[f(x)=-2x^2+6x\]

OpenStudy (anonymous):

b isn't -6?

OpenStudy (anonymous):

jsut kidding

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

lol thank you so much!

OpenStudy (anonymous):

you are a superhero

OpenStudy (anonymous):

spiderman?

OpenStudy (anonymous):

http://www.unc.edu/depts/jomc/academics/dri/idog.jpg

OpenStudy (anonymous):

nahh captain america! more hardcore

OpenStudy (anonymous):

ahahahahah